Route Canal Shipping Company has the following schedule for aging of accounts receivable: Age of Receivables April 30, 20X1 (1) (2) (3) (4) Month of Sales Age of Account Amounts Percent of Amount Due April 0–30 $ 131,760 _______ March 31–60 49,410 _______ February 61–90 115,290 _______ January 91–120 32,940 _______ Total receivables $ 329,400 100% a. Calculate the percentage of amount due for each month.

a. Calculate the percentage of amount due for each month.

b. If the firm had $1,464,000 in credit sales over the four-month period, compute the average collection period. Average daily credit sales should be based on a 120-day period.

c. If the firm likes to see its bills collected in 32 days, should it be satisfied with the average collection period?

d. Disregarding your answer to part c and considering the aging schedule for accounts receivable, should the company be satisfied?
Answer #1

a.

Month of sales Age of account Amount Percentage of amount due
April 0-30 131760 40%
March 31-60 49410 15%
February 61-90 115290 35%
January 91-120 32940 10%
329400

b. Average daily credit sales = 1,464,000 / 120 = 12200

Average collection period = 329400 / 12200 = 27 days

c. Yes

d. No .